# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ E.3 Problem State $C_{t}$
For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder.
For OP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100). The setting is the same as AM (Kool et al., 2019).
For PCTSP and SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al. (2019).
For the problem state, we provide an ablation study in Appendix K.2.
@@ E.4 Training
$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\bm{\theta}}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (16)
@@ Appendix F Adaptation Attention Free Module
Following Zhou et al. (2024a), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations:
@@
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. The AAFM computation is then expressed as:
@@
where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha$ is generated via a lightweight network $\mathrm{BIAS}(\bm{\lambda})$ in our paper (see Equation 6). Notably, for the relation matrix, we remove the scale $N$ in the calculation of adaptation bias because the relation is scale-independent.
Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work section mentioned above.
Appendix G LLM-driven Constraint Satisfaction Mechanism
To operationalize the UDR framework for a vast spectrum of routing problems, we introduce an automated constraint satisfaction mechanism based on LLMs. This mechanism obviates the need for manual, problem-specific masks by automatically designing executable mask generators. This approach follows the generate-and-validate pipeline, transforming constraint specifications into reliable code.
G.1 Mask Generator Design
This process begins by prompting an LLM with a prompt that comprises two key elements: 1) a natural language task description outlining the constraint logic, and 2) a standardized Python code template that specifies the function signature, inputs, and outputs. Here is an example for generating a mask for the TSP that enforces the fundamental visit-once constraint:
Task Description: Given the current node, visited nodes, and the number of nodes, design a function to mask visited nodes.
Code Template:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 return mask
The LLM completes the function body with the prompt guidance. The candidate mask generator is an executable Python snippet implementing the described logic:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 mask = torch.ones(current_node.size(0), num_nodes, dtype=torch.bool)
17 mask.scatter_(1, visited_nodes, 0)
18 mask.scatter_(1, current_node, 0)
19 return mask
G.2 Automated Validation
Each candidate generator is then validated. We execute it within the URS solver or a simple simulated solver on a set of problem instances and use a problem-specific checker function to verify if the resulting solutions are constraint-compliant. For the TSP example, the checker ensures that every node is visited exactly once:
1import torch
2
3def checker(solutions: torch.Tensor, problem_size: int, num_instances: int) -> float:
4 """
5 Check the validity of TSP solutions that use the generated masks.
6
7 Args:
8 solutions: A batch of generated solutions to be validated. Each row represents a single solution sequence. Shape: [batch_size, problem_size].
9 problem_size: The problem size.
10
11 Returns:
12 valid_rate: The validity rate, representing the fraction of valid solutions in the batch.
13 """
14 num_instances = solutions.shape[0]
15 sorted, _ = torch.sort(solutions, dim=1)
16 valid = (sorted == torch.arange(problem_size)[None, :].expand(num_instances, problem_size)).all(1)
17 valid_rate = torch.mean(valid.float()).item()
18 return valid_rate
G.3 AHD Framework Integration
The mask generator design process can be formulated as an optimization problem aimed at maximizing the solution validity rate across test instances. This allows us to leverage sophisticated LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), to improve the efficiency, especially for the complex VRP variants. The search terminates as soon as a fully reliable generator is found, making the process significantly more efficient than general optimization tasks. This LLM-driven synthesis, combined with UDR, empowers URS with true zero-shot generalization capabilities while rigorously upholding solution feasibility across an unprecedented range of routing problems.
@@ Appendix H Training and Model Settings
The detailed information about the hyperparameter settings can be found in Table 9. The training of URS is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory). The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
@@
Appendix I Detailed Experimental Results
To comprehensively evaluate zero-shot generalization of URS, we test its performance on a diverse set of 96 unseen problems. This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants from the NCO literature and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 10 and Table 11. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants, which compound asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for them, strengthening the broad problem generality and deployment efficiency.
Table 10: Zero-shot generalization performance across $44$ VRP variants that have been explored in current literature (Liu et al., 2024a; Berto et al., 2024) ($N=100$).
@@ Appendix I Detailed Experimental Results
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
@@
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
MVMoE 10.85% 15s 16.88% 14s 1.84% 16s 0.26% 12s 1.35% 12s
MTPOMO 11.53% 9s 16.39% 8s 2.77% 10s 0.48% 6s 1.79% 6s
[…]
Method OCVRPB CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h
MVMoE 7.09% 12s 1.47% 16s 9.95% 15s 7.33% 16s 3.15% 13s
MTPOMO 7.34% 5s 1.92% 9s 10.45% 6s 7.75% 8s 3.44% 6s
[…]
URS 9.35% 6s 2.67% 9s 13.77% 7s 9.18% 8s 3.22% 7s
@@
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 64.91% 33s
MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.66% 24s
[…]
@@
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 52.16% 33s 64.79% 36s 46.18% 38s 54.23% 42s 54.76% 39s
MTPOMO 40.15% 24s 45.84% 26s 34.19% 29s 40.26% 32s 40.32% 30s
ReLD-MTL 34.53% 27s 39.06% 29s 34.08% 34s 36.59% 37s 37.16% 35s
[…]
@@
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 51.72% 34s 45.23% 40s 31.90% 31s 53.64% 41s 29.08% 31s
MTPOMO 39.34% 27s 33.37% 32s 23.75% 23s 36.10% 31s 27.46% 23s
[…]
@@
AM - - - 2.69% 5s -
[…]
Table 11: Zero-shot generalization performance across $52$ VRP variants that have not yet been explored in the current literature ($N=100$).
@@
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method APDCVRP AOPDCVRP - - -
Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m - - -
[…]
Appendix J Results on Benchmark Dataset
@@
Our test results on Benchmark datasets are presented in Table 12. All models are trained on instances of size $N=100$. Some results for comparative models are sourced from MVMoE (Zhou et al., 2024b) and RouteFinder (Berto et al., 2024). URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based representative specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
@@
[…]
To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$. Table 13 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup–delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance. Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.

### v2

@@ E.3 Problem State $C_{t}$
For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder. For OP and AOP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100 and 1.0 in AOP100). The setting is the same as AM (Kool et al., 2019). For (A)PCTSP and (A)SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al. (2019). For the problem state, we provide an ablation study in Appendix K.2.
@@ E.4 Training
$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\theta}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (17)
@@ Appendix F Adaptation Attention Free Module
Following Zhou et al. (2026), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, the AAFM computation is then expressed as:
@@
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha$ is generated via a lightweight network $\mathrm{BIAS}(\bm{\lambda})$ in our paper (see Equation 6). Notably, for the relation matrix, we remove the scale $N$ in the calculation of adaptation bias because the relation is scale-independent. Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work mentioned above.
Appendix G Comparison between Different Attention Mechanisms
Table 13: Comparison between different attention mechanisms. Here, $N,d,m$ denote the sequence length, feature dimension, and extra hidden layer dimension of mixed-score attention, respectively. Values in parentheses represent the percentage increase or decrease relative to URS and Avg. Peak VRAM represents the average peak VRAM per instance.
Attention Mechanism Time Complexity Space Complexity #Params Avg.Peak VRAM Inference Time
Multi-Head Attention $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 3.323.32M(×0.7)($\times$ 0.7) 1.871.87MB(×0.99)($\times$ 0.99) $0.92$s(×0.9)($\times$ 0.9)
Mixed-Score Attention (Parallel) $\mathcal{O}(N^{2}(d+m))$ $\mathcal{O}(N^{2}m+Nd)$ 16.1716.17M(×3.3)($\times$ 3.3) 25.6425.64MB(×13.6)($\times$ 13.6) $3.64$s(×3.7)($\times$ 3.7)
Mixed Bias Module (Ours) $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 4.964.96M 1.891.89MB $0.97s$
To validate the efficiency of our proposed MBM, we have conducted a new comparative experiment between different attentions and tested them on TSP instances to ensure consistent decoding steps. Specifically, we replace our MBM with both a vanilla MHA (Vaswani et al., 2017) and the parallel mixed-score attention (MSA) used in MatNet (Kwon et al., 2021), and compare their complexities, memory overheads, and inference time. The results in Table 13 show that the increased VRAM and time overhead of MBM compared to standard MHA are small, and both maintain the same time complexity and space complexity. Compared to MatNet’s parallel MSA, our MBM drastically reduces both VRAM consumption and inference latency. These results show that MBM imposes only modest computational penalties while successfully capturing crucial problem-specific biases, which enhances cross-problem generalization while reducing architectural redundancy.
@@ Appendix H Training and Model Settings
H.1 Multi-Task Version
@@
The detailed information about the hyperparameter settings can be found in Table 14. The URS training is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB of memory). The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
H.2 Single-Task Version
URS-STL is trained on data from only a single problem type (e.g., ”URS-STL (CVRP)” is trained only on CVRP). The URS-STL variant uses exactly the same model architecture, hyperparameters (e.g., embedding dimension, number of layers, optimizer), and training setup (e.g., initial learning rate, batch size) as the multi-task URS described in Section 4.1 and Appendix H.1. The only modification we made is to the training duration. We observe that the single-task models converge significantly faster than the multi-task model. Therefore, we train URS-STL models for 300 epochs, with a learning rate decay applied at the 251st epoch. This contrasts with the multi-task URS, which is trained for 500 epochs with decay applied at the 451st epoch. This adjustment is made solely to reflect the faster convergence of the single-task training. All other hyperparameters and model architecture remain identical.
Appendix I Generalization to Unseen VRPs
To comprehensively evaluate URS’s zero-shot generalization, we test its performance on a diverse set of 99 unseen problems. This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants and 55 less-explored VRP variants. The detailed results are presented in Table 15 and Table 16. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations. Without any fine-tuning, URS still produces high-quality solutions for them, thereby strengthening its broad problem generality and deployment efficiency.
Table 15: Zero-shot generalization performance across $44$ unseen well-studied VRP variants ($N=100$).
@@ Appendix I Generalization to Unseen VRPs
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 18.27% 14s 13.38% 14s 9.42% 15s 16.66% 16s
MTPOMO 13.71% 7s 17.90% 8s 13.44% 8s 9.94% 9s 17.15% 10s
ReLD-MTL 13.57% 9s 15.81% 9s 12.96% 10s 8.43% 11s 15.95% 12s
URS 12.95% 7s 16.96% 7s 12.65% 8s 10.05% 8s 18.33% 10s
@@
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
MVMoE 16.86% 15s 17.75% 14s 8.87% 16s 0.26% 12s 1.35% 12s
MTPOMO 17.22% 9s 17.31% 8s 9.47% 10s 0.48% 6s 1.79% 6s
ReLD-MTL 16.09% 11s 15.21% 10s 8.06% 11s 0.02% 8s 1.01% 7s
URS 18.15% 9s 16.37% 8s 9.51% 9s 0.43% 7s 1.65% 6s
@@
Oracle 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 65.98% 33s
MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.77% 24s
ReLD-MTL 6.74% 8s 5.41% 7s 3.16% 9s 9.22% 8s 40.70% 27s
URS 8.94% 8s 9.47% 7s 5.12% 9s 13.72% 8s 35.44% 24s
@@
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 56.60% 33s 65.81% 36s 63.77% 38s 68.49% 42s 69.00% 39s
MTPOMO 42.94% 24s 45.98% 26s 47.09% 29s 50.45% 32s 50.74% 30s
ReLD-MTL 37.74% 27s 39.57% 29s 45.16% 34s 45.23% 37s 45.62% 35s
URS 24.44% 22s 35.07% 26s 26.31% 26s 36.49% 32s 36.71% 30s
@@
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 56.26% 34s 62.47% 40s 33.06% 31s 53.64% 41s 29.08% 31s
MTPOMO 41.98% 27s 46.26% 32s 23.38% 23s 36.10% 31s 27.46% 23s
ReLD-MTL 37.45% 29s 44.74% 36s 18.39% 25s 30.90% 34s 22.32% 25s
URS 24.22% 24s 26.20% 28s 15.05% 23s 37.26% 32s 22.01% 23s
@@
URS 14.90% 20s 25.59% 27s 30.41% 24s -2.37% 5s -
Table 16: Zero-shot generalization performance across $55$ unseen less-explored VRP variants ($N=100$).
@@
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 39.31% 6.7m 24.99% 6.9m 20.98% 7.6m -1.47% 4.3s 4.93% 4.3s
Method APDCVRP AOPDCVRP AOP APCTSP ASPCTSP
Gap Time Gap Time Gap Time Gap Time Value Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m -
URS 7.03% 1.2m 11.13% 1.2m -5.47% 1.2m 43.79% 1.2m 2.22 1.2m
Appendix J More Experiments
J.1 Training Problem Selection
To empirically validate our selection strategy of training problems, we train two additional variants and compare them with our original URS: URS-Diff (Different Selection) and URS-Expand (Broader Problems). The first variant, URS-Diff, uses a different problem set where OCVRPTW is replaced by attribute-independent CVRPL. This modification means that the Open Route and Time Window attributes each appear in only one training problem. The second variant, URS-Expand, is trained on a broader set of 22 problems, incorporating all variants evaluated in MTPOMO (Liu et al., 2024) and MVMoE (Zhou et al., 2024).
As shown in Table 17, URS-Diff performs significantly worse on unseen problems involving Open Route and Time Window constraints. Meanwhile, URS-Expand achieves a slightly better performance than URS due to seeing more problem variants during training. However, the original URS remains highly competitive. Remarkably, despite being trained on only half the number of problem types (11 vs. 22), URS achieves the best solution on 12 out of the 22 evaluated problems.
These results demonstrate that the choice of our selected 11 problems offers a balanced trade-off, providing sufficient attribute coverage for strong generalization while maintaining training efficiency. Expanding the training set yields only marginal gains, while a poorly selected set leads to poor cross-problem generalization. In future work, we plan to explore more effective training problem selection strategies to further enhance model performance and training efficiency.
Table 17: Comparison under different training problem settings. The symbol * indicates the original problem selection, and $\Delta$ represents a different problem selection of equal number. All test problems are encompassed within URS-Expand.
Method ATSP (* $\Delta$) TSP(* $\Delta$) OP(* $\Delta$) PCTSP(* $\Delta$) PDTSP(* $\Delta$) ACVRP($\Delta$)
URS-Diff 4.52% 1.05% 0.82% 1.31% 6.09% 5.10%
URS-Expand 4.77% 0.93% 0.99% 1.41% 5.85% 5.48%
URS 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
CVRP(* $\Delta$) CVRPTW(* $\Delta$) CVRPB(* $\Delta$) OCVRP(* $\Delta$) OCVRPTW(*) CVRPL($\Delta$)
URS-Diff 1.97% 6.26% 1.80% 3.75% 15.45% 0.58%
URS-Expand 2.04% 5.62% 1.77% 3.45% 4.34% 0.65%
URS 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
URS-Diff 7.55% 2.02% 2.76% 27.48% 10.25% 3.73%
URS-Expand 5.39% 1.88% 2.03% 9.13% 6.29% 3.41%
URS 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
URS-Diff 10.00% 7.56% 15.61% 27.80% 7.43% 0/22
URS-Expand 6.07% 5.47% 4.38% 9.23% 4.12% 10/22
URS 8.94% 9.47% 5.12% 13.72% 4.89% 12/22
J.2 Comparison with RouteFinder and CaDA
We strictly adhered to the problem settings established by MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025) to ensure a fair and consistent evaluation. Notably, as shown in Table 6, URS already outperforms RF (Berto et al., 2025) and CaDA (Li et al., 2025a) on the large-scale CVRPLib benchmark dataset. In our cross-problem test, RF and CaDA are excluded due to discrepancies in the problems, including the number of trained CVRP variants (16 in their work vs. 5 in ours) and constraint settings (i.e., Backhauls and Time Windows). To ensure a fair comparison with RF and CaDA, we conduct two additional experiments:
  • We evaluate URS against CaDA and the best-performing RouteFinder variant (RF-TE) on a shared subset of both seen and unseen problems, strictly restricting the evaluation to variants with perfectly aligned constraint definitions across all methods. As shown in Table 18, when evaluated under strictly matched conditions, URS still performs best overall among comparable methods. Note that URS can provide the strongest zero-shot generalization across all shared unseen problems.
  • We retrain both RouteFinder-TE (denoted as RF-Retrain) and CaDA (denoted as CaDA-Retrain) from scratch using the exact same training settings as our URS and evaluate the three models on 48 CVRP variants. Since RF fails to train beyond 150 epochs due to the occurrence of NaN probabilities in our training setting, we evaluate all three methods at exactly 150 epochs to ensure a fair comparison. The experimental results in Table 19 show that under consistent training settings, URS still achieves the best overall performance among the three methods and provides the best solutions for the majority of problems (32/48), highlighting our strong zero-shot generalization.
Table 18: Comparison of shared seen and unseen VRP variants between URS, RF-TE, and CaDA.
Method Seen Unseen Avg.gap Best Sol.
CVRP OCVRP MDCVRP MDOCVRP MDCVRPL MDOCVRPL
RF-TE 2.06% 2.86% 63.60% 33.52% 66.75% 34.40% 33.86% 0/6
CaDA 2.17% 2.80% 40.96% 53.21% 42.40% 53.34% 32.48% 1/6
URS 1.81% 3.24% 15.04% 22.01% 15.32% 22.15% 13.26% 5/6
Table 19: Comparison between URS and the retrained RouteFinder and CaDA on all 48 CVRP variants. Note that all models are trained for 150 epochs to ensure a fair comparison.
Method CVRPTW OCVRP CVRP CVRPB OCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 8.38% 8s 6.03% 6s 3.58% 6s 4.47% 5s 8.18% 7s
CaDA-Retrain 5.28% 9s 3.22% 7s 1.92% 7s 1.61% 6s 4.42% 8s
URS 6.09% 8s 3.71% 6s 2.07% 6s 1.90% 5s 4.89% 8s
Method OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 11.61% 5s 4.73% 6s 4.77% 9s 14.17% 7s 10.11% 8s
CaDA-Retrain 8.46% 6s 4.42% 7s 4.63% 11s 10.39% 8s 10.87% 10s
URS 8.01% 6s 2.12% 6s 2.53% 9s 12.37% 7s 8.74% 8s
Method OCVRPL CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 5.98% 7s 9.74% 7s 11.58% 6s 8.34% 8s 14.46% 7s
CaDA-Retrain 5.81% 8s 7.17% 8s 12.48% 7s 8.55% 11s 15.94% 10s
URS 3.72% 7s 8.42% 7s 8.05% 6s 5.06% 8s 12.37% 8s
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 14.40% 7s 18.31% 7s 14.01% 7s 11.43% 9s 18.70% 10s
CaDA-Retrain 13.37% 8s 18.25% 8s 16.71% 9s 9.64% 10s 20.17% 12s
URS 13.37% 7s 16.08% 7s 12.94% 8s 10.07% 9s 18.47% 10s
Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL MDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 18.84% 9s 17.76% 8s 10.88% 9s 2.21% 7s 27.10% 18s
CaDA-Retrain 17.11% 10s 21.44% 9s 13.54% 12s 2.29% 8s 27.09% 23s
URS 18.63% 9s 15.42% 8s 9.57% 10s 0.66% 7s 38.17% 23s
Method MDCVRPTW MDOCVRP MDCVRPL MDCVRPB MDOCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 76.46% 25s 19.09% 18s 27.41% 34s 21.40% 15s 35.23% 22s
CaDA-Retrain 51.08% 30s 30.01% 19s 27.59% 21s 16.47% 19s 54.46% 29s
URS 38.25% 31s 21.34% 21s 39.22% 26s 20.83% 18s 28.40% 22s
Method MDOCVRPB MDCVRPBL MDCVRPLTW MDOCVRPBTW MDCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 17.84% 15s 20.02% 29s 79.33% 40s 44.11% 19s 82.19% 23s
CaDA-Retrain 26.72% 16s 15.02% 18s 43.43% 29s 64.29% 25s 50.11% 24s
URS 15.20% 17s 17.36% 21s 39.98% 33s 25.68% 19s 33.24% 27s
Method MDOCVRPL MDCVRPBTW MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 19.02% 19s 84.94% 21s 17.72% 16s 35.06% 24s 44.12% 21s
CaDA-Retrain 30.61% 20s 61.39% 25s 26.94% 17s 39.49% 29s 49.12% 25s
URS 21.56% 24s 33.58% 25s 14.84% 19s 28.60% 24s 25.84% 21s
Method MDCVRPBP MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 54.74% 21s 36.08% 21s 54.19% 23s 47.24% 26s 84.51% 30s
CaDA-Retrain 54.45% 25s 52.38% 22s 50.69% 24s 69.07% 32s 55.63% 32s
URS 41.49% 24s 24.18% 23s 41.08% 27s 24.19% 25s 33.99% 35s
Method MDCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW Avg.gap Best Solutions
Gap Time Gap Time Gap Time
RF-Retrain 85.12% 28s 36.06% 23s 47.63% 28s 29.24% 3/48
CaDA-Retrain 66.78% 33s 53.48% 23s 53.06% 31s 28.45% 13/48
URS 34.55% 32s 24.10% 25s 24.32% 27s 19.01% 32/48
J.3 Detailed Results on CVRPLib Set-X Dataset
@@
We further evaluate the generalization ability of URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$. Our test results on CVRPLib Set-X datasets are presented in Table 20. All models are trained on instances of size $N=100$. In the Set-X dataset, some results for comparative models are sourced from MVMoE (Zhou et al., 2024) and RouteFinder (Berto et al., 2025). As shown in Table 20, URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based representative specialist model LEHD (8.678% vs. 12.836%) and many representative multi-task neural solvers (e.g., MVMoE and RouteFinder), underscoring its practical applicability in real-world scenarios.
@@
w/o $\bm{\xi}_{i}$ 2.14% 0.48% 0.45% 1.06% 9.28% 2.84% 3.66% 6.61% 1.57% 7.87% 5.66% 3.78%
w/ $\bm{\xi}_{i}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$. Table 21 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup-delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance. Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.